# Fractional Sobolev paths on Wasserstein spaces and their energy-minimizing particle representations

### v3

@@ 1.2. Motivating example
In the first case of this example, we revisit the well-known relationship between the continuity equation and absolutely continuous Wasserstein curves under the $p$-integrability condition, [4, Chapter 8] and [14, Section 1]. In the second case, we assume a weaker integrability condition and highlight the estimates applied in the fractional setting.
@@
with initial condition $\mu_{0}\in P(\mathbb{R}^{\mathrm{d}})$ and for some Borel velocity vector field $v:[0,T]\times\mathbb{R}^{\mathrm{d}}\to\mathbb{R}^{\mathrm{d}}$, where $v_{t}(x)\coloneqq v(t,x)$. The interpretation of this equation in the distributional sense (i.e. in duality with smooth functions with compact support), which we also consider here, makes sense without any regularity assumption on $v$ provided an integrability condition for the couple $(\mu_{t},v_{t})$ such as the (global) 1-integrability condition:
@@
which also guarantees the conservation of mass. When, in addition, $v$ is regular enough (e.g. its Lipschitz and supremum norms over compact sets are time-integrable [4, Eq. (8.1.8)]), the associated ODE
$\begin{cases}\dot{X}_{t}(x)&=v_{t}(X_{t}(x))\qquad\text{in }[0,T],\\ X_{0}(x)&=x,\end{cases}$ (1.23)
admits a unique global solution. In this case, we know (e.g. by [4, Proposition 8.1.8]) that $(\mu_{t})$ is represented by
$\mu_{t}=(X_{t})_{\#}\mu_{0}\qquad\forall t\in[0,T],$ (1.24)
where $X_{t}:x\mapsto X_{t}(x)$ is a map from $\mathbb{R}^{\mathrm{d}}\to\mathbb{R}^{\mathrm{d}}$. In this case, $(\mu_{t})$ has a lift $\pi\in P(\Gamma_{T})$ given by
$\pi\coloneqq(X)_{\#}\mu_{0}$ (1.25)
where $X:x\mapsto(X_{t}(x))_{t\in I}$ is a map from $\mathbb{R}^{\mathrm{d}}\to\Gamma_{T}$, mapping the initial datum to the entire solution curve. From now on in this example, we assume that the 1-integrability condition (1.22) and the regularity assumption is in force, ensuring the existence of a lift $\pi$.
@@
for some $1<p<\infty$. The first implication of these conditions is that they ensure $\mu_{t}\in P_{p}(\mathbb{R}^{\mathrm{d}})$ for all $t\in(0,T]$. We can rewrite the expression above as the $W^{1,p}$-energy of the lift (1.25):
@@
where $W^{1,p}$-regularity of an $p$-absolutely continuous curve $\gamma:[0,T]\to\mathbb{R}^{\mathrm{d}}$ over an interval $[s,t]\subset[0,T]$ is given by (see Section 2.5 for the definition of $W^{1,p}$-regularity)
@@
and we shall omit the time interval when $[s,t]=[0,T]$. A simple application of Hölder’s inequality yields
@@
This can be used to infer that $(\mu_{t})$ is an absolutely continuous curve. Moreover, as shown in [14, Theorem 4] by the Lebesgue differentiation theorem, we obtain an upper bound for the metric speed $|\dot{\mu}_{t}|\leq\|\dot{\gamma}_{t}\|_{L^{p}(\pi)}$ for a.e. $t\in I$. Raising it to power $p$ and integrating in time yields
@@
Case 2 ($W^{\alpha,p}$-regularity). We shall still assume $v$ to be regular enough and satisfy the 1-integrability condition (1.22), but not the $p$-integrability or its equivalent formulation (1.27) (we also assume that there is no other velocity vector field satisfying the $p$-integrability and the continuity equation). Instead, let us assume a weaker integrability condition
$\int_{\Gamma_{T}}|\gamma|^{p}_{W^{\alpha,p};[0,T]}{\mathrm{\,d}}\pi(\gamma)<+\infty$ (1.30)
for some $1<p<\infty$ and $11/p<\alpha<$ 1, but with the same initial condition $\mu_{0}\in P_{p}(\mathbb{R}^{\mathrm{d}})$. These again ensure $(\mu_{t})\subset P_{p}(\mathbb{R}^{\mathrm{d}})$ by Theorem 1.2. Such a situation as above is present in Example 4.3, where the superposition of absolutely continuous curves results in, instead of an absolutely continuous Wasserstein curve, a Hölder curve of infinite length. To proceed, let us denote the $W^{\alpha,p}$-regularity of a curve $\gamma:[0,T]\to\mathbb{R}^{\mathrm{d}}$ over an interval $[s,t]\subset[0,T]$ by
@@
In the fractional setting, the Hölder’s inequality (1.28) is replaced by the Garsia–Rodemich–Rumsey inequality [11, 10]
@@ 1.4. Acknowledgements
The author would like to thank his supervisor, Matthias Erbar, for invaluable guidance and numerous discussions. He also thanks Timo Schultz for insightful conversations, which significantly improved the results and eventually led to Examples 4.3, 4.4 and 4.5. He also thanks Zhenhao Li for helpful comments on some technical details. Additionally, he extends his thanks to Kohei Suzuki and Vitalii Konarovskyi for insightful discussions. The author acknowledges the financial support of the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.
@@ 2.1. Continuous paths on metric spaces
Throughout the paper, $(\mathcal{X},d)$ is a metric space and $[0,T]\subset\mathbb{R}$ is a time interval. Any additional structure on $(\mathcal{X},d)$ will be explicitly stated. Let $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ denote the space of all continuous paths $X:[0,T]\to\mathcal{X}$. When $T=1$, we write $\Gamma\coloneqq\Gamma_{1}$. For any two paths $X,Y\in\Gamma_{T}$, the supremum distance is defined by
@@
The Arzelà-Ascoli theorem provides a characterization of relative compactness in $\Gamma_{T}$ (see, e.g., [6, Theorem 7.2] or [10, Theorem 1.4]).
@@ 2.2. Hölder paths on metric spaces
Given $\upgamma\in[0,1]$, the $\upgamma$-Hölder continuity of a path $X\in C([0,T];\mathcal{X})$ over $[s,t]\subset[0,T]$ is defined by
@@
Below, a simple yet useful characterization of Hölder continuity is provided. It states that for a continuous curve $X:[0,1]\to\mathcal{X}$ to be Hölder, it is enough to verify the Hölder condition only at the dyadic time points of $[0,1]$. The result is taken from [18, Lemma 2], whose proof is inspired by the classical proof of Kolmogorov–Čentsov continuity theorem (see e.g. [13, Theorem 2.8]).
Theorem 2.7 (A discrete characterization of $C^{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$ [18, Lemma 2]).
@@
We recall some properties of the Hölder regularity. For below, see e.g. [10, Lemma 5.12].
@@ 2.3. $p$-variation paths on metric spaces
be a dissection of the time interval $[s,t]\subset\mathbb{R}$. The mesh of $D$ is defined as
@@
Given $p\geq 1$, the $p$-variation of a path $X\in C([0,T];\mathcal{X})$ over $[s,t]\subset[0,T]$ is defined by
@@
Two subsequent lemmas are taken from [10, Proposition 5.6].
@@
To distinguish these norms, we refer to the first one as variation, while the second one as infinitesimal variation, emphasizing that they are not necessarily identical. This is only the case for $p=1$ (see e.g. [10, Proposition 1.14]):
@@
As a next result, we recall the lower semi-continuity of $p$-variation, as stated in [10, Lemma 5.12].
@@
Let $(\mathcal{X},d)$ be a geodesic space. Take $p>1$ and let $\gamma$ be a continuous path such that $0<|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}<\infty$. Let $(D_{n})_{n\in\mathbb{N}}$ be a sequence of dissections of $[0,T]$ with shrinking mesh size $|D_{n}|\to 0$. Define $\gamma^{D_{n}}$ to be the piecewise geodesic approximation of $\gamma$ (i.e. the curve that coincides with $\gamma$ at points of $D_{n}$ and connects in between with geodesics). We know that $\gamma^{D_{n}}\to\gamma$ uniformly (see e.g. [10, Lemma 5.19]). For each $n$, the path $\gamma^{D_{n}}$ is of bounded $1$-variation, which implies, by Lemma 2.12, that $|\gamma^{D_{n}}|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}=0$ for all $p>1$. Consequently,
@@ 2.4. Fractional Sobolev paths on metric spaces
Given $1\leq p<\infty$ and $10<\alpha<$ 1, the fractional Sobolev regularity of a measurable function $X:[0,T]\to\mathcal{X}$ over $[s,t]\subset[0,T]$ is defined by
@@
The fractional Sobolev spaces $W^{\alpha,p}$ (also known as Sobolev–Slobodeckij spaces) can be viewed as a specific instance of the broader class of Besov spaces $B^{\alpha,p,q}$ (see e.g. [16, Equation (2.1)]). In particular, when $q=p$, we have $B^{\alpha,p,p}=W^{\alpha,p}$. While the classical definition of $B^{\alpha,p,q}$ typically involves double integrals similar to the one mentioned above, there are also other function spaces whose definition involves double sums. Recently, [16] has proved that these definitions are in fact equivalent. For the purpose of our work, we only give their result for the case $q=p$:
Theorem 2.17 (A discrete characterization of $W^{\alpha,p}$ [16, Theorem 2.2]).
@@
This result is reminiscent of the discrete characterization of Hölder curves provided by [18, Lemma 2], as stated in Theorem 2.7. In the second definition, we are once again evaluating $X$ only at dyadic points (i.e. a countable subset of $[0,1]$) but this time, in the form of a double sum. Thus, continuity must be an additional assumption. This is, however, not needed for the first definition. In fact, under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the finiteness of $|X|_{W^{\alpha,p}}$ automatically implies the continuity of $X$ (see Theorem 2.19). We repeatedly rely on the equivalence result above. We proceed with the following simple observation:
@@
The estimate (2.29) below can be derived by applying the Garsia–Rodemich–Rumsey inequality, which was originally introduced in [11] for the case $\mathcal{X}=\mathbb{R}$. A proof for $\mathbb{R}^{\mathrm{d}}$ can be found in [22, Theorem 2.1.3] and for the general metric setting in [10, Theorem A.1]). The results below are well-known [10, Corollary A.2-3].
@@
Note that from (2.29), we can immediately conclude that $X$ has finite $({\alpha-1/p})^{-1}$-variation. But, as shown in [9, Theorem 2], by applying Hölder’s inequality with a clever choice of exponents, an even stronger statement can be made, namely, $X$ has finite $({\alpha})^{-1}$-variation. A consequence of this embedding is that under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the elements in $W^{\alpha,p}$ have zero infinitesimal $p$-variation by Lemma 2.12:
@@
where we used Fatou’s lemma and (2.36). The final claim of the proposition follows from the fact that uniform convergence implies pointwise convergence. ∎
@@ 2.5. Sobolev paths on metric spaces
We know that the metric derivative $|\dot{X}_{t}|$ of any $X\in AC^{1}([0,T];\mathcal{X})$ exists a.e. $t\in[0,T]$ [4]. For any $X\in C([0,T];\mathcal{X})$ and any $[s,t]\subset[0,T]$, we set:
@@
Given $1<p<\infty$, the Sobolev regularity of a function $X:[0,T]\to\mathcal{X}$ over $[t,s]\subset[0,T]$ is defined by
@@
On $\mathbb{R}^{\mathrm{\,d}}$, it is well-known that the function spaces mentioned above coincide when $p>1$ (see e.g. [10, Proposition 1.45]). This result also holds in the general metric setting. For a proof in the case $p=2$, see [2, Theorem 10.2], which can be generalized to any $p>1$ and is stated below:
@@
Accordingly, we mainly focus on the definition $|\cdot|_{W^{1,p}}$ as its properties are easier to observe. First, since it is defined as the supremum of a family of continuous functions, it is a lower semi-continuous map, as mentioned in [2, Theorem 10.2]:
@@
which implies the embeddings [10, Theorem 1.47]

### v4

@@ 1.2. Motivating example
In the first case of this example, we revisit the well-known relationship between the continuity equation and absolutely continuous Wasserstein curves under the $p$-integrability condition, [6, Chapter 8] and [17, Section 1]. In the second case, we assume a weaker integrability condition and highlight the estimates applied in the fractional setting.
@@
with initial condition $\mu_{0}\in P(\mathbb{R}^{\mathrm{d}})$ and for some Borel velocity vector field $v:[0,T]\times\mathbb{R}^{\mathrm{d}}\to\mathbb{R}^{\mathrm{d}}$, where $v_{t}(x)\coloneqq v(t,x)$. Throughout this example, we assume that the couple $(\mu_{t},v_{t})$ satisfies the (global) 1-integrability condition:
@@
which guarantees the conservation of mass. We interpret (1.21) in the distributional sense (i.e., in duality with smooth functions with compact support). Under the assumption (1.22), the superposition principle of [3] ensures the existence of a path measure $\pi\in P(C(I;\mathbb{R}^{\mathrm{d}}))$ such that
  1. (i)
    $\pi$ is concentrated on the set of curves $\gamma\in AC(I;\mathbb{R}^{\mathrm{d}})$ with $\dot{\gamma}_{t}=v_{t}(\gamma_{t})$ for a.e. $t\in I$;
  2. (ii)
    $(e_{t})_{\#}\pi=\mu_{t}$ for all $t\in I$.
@@
for some $1<p<\infty$. The first implication of these conditions is that they ensure $\mu_{t}\in P_{p}(\mathbb{R}^{\mathrm{d}})$ for all $t\in(0,T]$. We can rewrite (1.23) as the $W^{1,p}$-energy of the lift $\pi$:
@@
where $W^{1,p}$-regularity of an $p$-absolutely continuous curve $\gamma:[0,T]\to\mathbb{R}^{\mathrm{d}}$ over an interval $s,t]\subset[0,T]$ is given by (see Section 2.5 for the definition of $W^{1,p}$-regularity)
@@
and we omit the time interval when $[s,t]=[0,T]$. An application of Hölder's inequality yields
@@
This can be used to infer that $(\mu_{t})$ is an absolutely continuous curve. Moreover, as shown in [17, Theorem 4] by the Lebesgue differentiation theorem, we obtain an upper bound for the metric speed $|\dot{\mu}_{t}|\leq\|\dot{\gamma}_{t}\|_{L^{p}(\pi)}$ for a.e. $t\in I$. Raising it to power $p$ and integrating in time yields
@@
Case 2 ($W^{\alpha,p}$-regularity). We still assume that $(\mu_{t},v_{t})$ satisfy the 1-integrability condition (1.22), but not the $p$-integrability or its equivalent formulation (1.24) for $p>1$ (we also assume that there is no other velocity vector field satisfying the $p$-integrability and the continuity equation). Instead, let us assume a weaker integrability condition directly on $\pi$:
$\int_{\Gamma_{T}}|\gamma|^{p}_{W^{\alpha,p};[0,T]}{\mathrm{\,d}}\pi(\gamma)<+\infty,$ (1.27)
for some $1<p<\infty$ and $1/p<\alpha<1$, with the same initial condition $\mu_{0}\in P_{p}(\mathbb{R}^{\mathrm{d}})$. These again ensure $(\mu_{t})\subset P_{p}(\mathbb{R}^{\mathrm{d}})$ by Theorem 1.2. Such a situation as above is present in Example 4.3, where the superposition of absolutely continuous curves results in, instead of an absolutely continuous Wasserstein curve, a Hölder curve of infinite length. To proceed, let us denote the $W^{\alpha,p}$-regularity of a curve $\gamma:[0,T]\to\mathbb{R}^{\mathrm{d}}$ over an interval $s,t]\subset[0,T]$ by
@@
In the fractional setting, the Hölder's inequality (1.25) is replaced by the Garsia–Rodemich–Rumsey inequality [14, 13]
@@ 1.4. Acknowledgements
The author would like to thank his supervisor, Matthias Erbar, for invaluable guidance and numerous discussions. He also thanks Timo Schultz for insightful conversations, which significantly improved the results and eventually led to Examples 4.3, 4.4 and 4.5. He also thanks Zhenhao Li for helpful comments on some technical details. Additionally, he extends his thanks to Kohei Suzuki and Vitalii Konarovskyi for insightful discussions. This work is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 317210226 – SFB 1283.
@@ 2.1. Continuous paths on metric spaces
Throughout the paper, $(\mathcal{X},d)$ is a metric space and $0,T]\subset\mathbb{R}$ is a time interval. Any additional structure on $(\mathcal{X},d)$ will be explicitly stated. Let $\Gamma_{T}\coloneqq C([0,T];\mathcal{X})$ denote the space of all continuous paths $X:[0,T]\to\mathcal{X}$. When $T=1$, we write $\Gamma\coloneqq\Gamma_{1}$. For any two paths $X,Y\in\Gamma_{T}$, the supremum distance is defined by
@@
The Arzelà-Ascoli theorem provides a characterization of relative compactness in $\Gamma_{T}$ (see, e.g., [8, Theorem 7.2] or [13, Theorem 1.4]).
@@ 2.2. Hölder paths on metric spaces
Given $\upgamma\in[0,1]$, the $\upgamma$-Hölder continuity of a path $X\in C([0,T];\mathcal{X})$ over $s,t]\subset[0,T]$ is defined by
@@
Below, a simple yet useful characterization of Hölder continuity is provided. It states that for a continuous curve $X:[0,1]\to\mathcal{X}$ to be Hölder, it is enough to verify the Hölder condition only at the dyadic time points of $[0,1]$. The result is taken from [21, Lemma 2], whose proof is inspired by the classical proof of Kolmogorov–Čentsov continuity theorem (see e.g. [16, Theorem 2.8]).
Theorem 2.7 (A discrete characterization of $C^{\upgamma\textrm{-}\mathrm{H\ddot{o}l}}$ [21, Lemma 2]).
@@
We recall some properties of the Hölder regularity. For below, see e.g. [13, Lemma 5.12].
@@ 2.3. $p$-variation paths on metric spaces
be a dissection of the time interval $s,t]\subset\mathbb{R}$. The mesh of $D$ is defined as
@@
Given $p\geq 1$, the $p$-variation of a path $X\in C([0,T];\mathcal{X})$ over $s,t]\subset[0,T]$ is defined by
@@
Two subsequent lemmas are taken from [13, Proposition 5.6].
@@
To distinguish these norms, we refer to the first one as variation, while the second one as infinitesimal variation, emphasizing that they are not necessarily identical. This is only the case for $p=1$ (see e.g. [13, Proposition 1.14]):
@@
As a next result, we recall the lower semi-continuity of $p$-variation, as stated in [13, Lemma 5.12].
@@
Let $(\mathcal{X},d)$ be a geodesic space. Take $p>1$ and let $\gamma$ be a continuous path such that $0<|\gamma|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}<\infty$. Let $(D_{n})_{n\in\mathbb{N}}$ be a sequence of dissections of $[0,T]$ with shrinking mesh size $|D_{n}|\to 0$. Define $\gamma^{D_{n}}$ to be the piecewise geodesic approximation of $\gamma$ (i.e. the curve that coincides with $\gamma$ at points of $D_{n}$ and connects in between with geodesics). We know that $\gamma^{D_{n}}\to\gamma$ uniformly (see e.g. [13, Lemma 5.19]). For each $n$, the path $\gamma^{D_{n}}$ is of bounded $1$-variation, which implies, by Lemma 2.12, that $|\gamma^{D_{n}}|_{p\textrm{-}\mathrm{var}\textrm{-}\mathrm{limsup}}=0$ for all $p>1$. Consequently,
@@ 2.4. Fractional Sobolev paths on metric spaces
Given $1\leq p<\infty$ and $0<\alpha<1$, the fractional Sobolev regularity of a measurable function $X:[0,T]\to\mathcal{X}$ over $s,t]\subset[0,T]$ is defined by
@@
The fractional Sobolev spaces $W^{\alpha,p}$ (also known as Sobolev–Slobodeckij spaces) can be viewed as a specific instance of the broader class of Besov spaces $B^{\alpha,p,q}$ (see e.g. [19, Equation (2.1)]). In particular, when $q=p$, we have $B^{\alpha,p,p}=W^{\alpha,p}$. While the classical definition of $B^{\alpha,p,q}$ typically involves double integrals similar to the one mentioned above, there are also other function spaces whose definition involves double sums. Recently, [19] has proved that these definitions are in fact equivalent. For the purpose of our work, we only give their result for the case $q=p$:
Theorem 2.17 (A discrete characterization of $W^{\alpha,p}$ [19, Theorem 2.2]).
@@
This result is reminiscent of the discrete characterization of Hölder curves provided by [21, Lemma 2], as stated in Theorem 2.7. In the second definition, we are once again evaluating $X$ only at dyadic points (i.e. a countable subset of $[0,1]$) but this time, in the form of a double sum. Thus, continuity must be an additional assumption. This is, however, not needed for the first definition. In fact, under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the finiteness of $|X|_{W^{\alpha,p}}$ automatically implies the continuity of $X$ (see Theorem 2.19). We repeatedly rely on the equivalence result above. We proceed with the following simple observation:
@@
The estimate (2.29) below can be derived by applying the Garsia–Rodemich–Rumsey inequality, which was originally introduced in [14] for the case $\mathcal{X}=\mathbb{R}$. A proof for $\mathbb{R}^{\mathrm{d}}$ can be found in [26, Theorem 2.1.3] and for the general metric setting in [13, Theorem A.1]). The results below are well-known [13, Corollary A.2-3].
@@
Note that from (2.29), we can immediately conclude that $X$ has finite $({\alpha-1/p})^{-1}$-variation. But, as shown in [12, Theorem 2], by applying Hölder's inequality with a clever choice of exponents, an even stronger statement can be made, namely, $X$ has finite $({\alpha})^{-1}$-variation. A consequence of this embedding is that under the conditions $1<p<\infty$ and $\frac{1}{p}<\alpha<1$, the elements in $W^{\alpha,p}$ have zero infinitesimal $p$-variation by Lemma 2.12:
@@
where we used Fatou's lemma and (2.36). The final claim of the proposition follows from the fact that uniform convergence implies pointwise convergence. ∎
@@ 2.5. Sobolev paths on metric spaces
We know that the metric derivative $|\dot{X}_{t}|$ of any $X\in AC^{1}([0,T];\mathcal{X})$ exists a.e. $t\in[0,T]$ [6]. For any $X\in C([0,T];\mathcal{X})$ and any $s,t]\subset[0,T]$, we set:
@@
Given $1<p<\infty$, the Sobolev regularity of a function $X:[0,T]\to\mathcal{X}$ over $t,s]\subset[0,T]$ is defined by
@@
On $\mathbb{R}^{\mathrm{\,d}}$, it is well-known that the function spaces mentioned above coincide when $p>1$ (see e.g. [13, Proposition 1.45]). This result also holds in the general metric setting. For a proof in the case $p=2$, see [4, Theorem 10.2] or [10, Proposition 2.3], which can be generalized to any $p>1$ and is stated below:
@@
Accordingly, we mainly focus on the definition $|\cdot|_{W^{1,p}}$ as its properties are easier to observe. First, since it is defined as the supremum of a family of continuous functions, it is a lower semi-continuous map, as mentioned in [4, Theorem 10.2]:
@@
which implies the embeddings [13, Theorem 1.47]